Kyky232:

(PIC IN REPLIES) segment AB is dilated from the origin to create segment A prime B prime at A' (0, 8) and B' (8, 12). What scale factor was segment AB dilated by? one half 2 3 4

Kyky232:

1 attachment
freshavacado850:

2

Kyky232:

just letting you know you were wrong

GalaxyzStarz:

abraham lincoln

freshavacado850:

lemme check again

freshavacado850:

b

Kyky232:

no...

Kyky232:

it was 1/2 I think

freshavacado850:

a=(0,8) b = (8,12 a=(0,4) b = 4,6 a = k(0,4)=(k,4k)=(0,8) k=0 4k=8 k=8/4 k=2

freshavacado850:

the value of AB=2

Kyky232:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @freshavacado850 the value of AB=2 \(\color{#0cbb34}{\text{End of Quote}}\) Well I put that and it was wrong. So no.